# High-dimensional quantum XYZ product codes for biased noise

### v2

@@ Abstract
Quantum XYZ product can construct a class of non-CSS codes by using three classical codes. However, before this work, their error-correcting performance is not studied in depth and whether this code construction method can be generalized to higher dimension is an open question. In this paper, we first study the error-correcting performance of the 3D Chamon code, which can be seen as a non-CSS variant of the 3D toric code and a special instance of the XYZ product of three repetition codes. Second, we show that XYZ product can be generalized to four dimension and propose four-dimensional (4D) XYZ product code construction, which can be seen as a variant of 4D homological product and constructs a class of non-CSS codes by using 4 classical codes or 2 CSS codes. Compared with 4D homological product, we show that 4D XYZ product can construct non-CSS codes with higher dimension or code distance. Third, we consider two special instances of 4D XYZ product, which we name 4D Chamon code and 4D XYZ concatenated code. Exploiting fully decoupled binary belief propagation combined with $order\mbox{-}0$ ordered statistics decoding, our simulation results show that, using the same two CSS codes, 4D XYZ product can construct non-CSS codes with better error-correcting performance for $Z$-biased noise than CSS codes constructed by 4D homological product, which is more meaningful for practice quantum computing system.
I Introduction
At present, even for the state-of-the-art quantum computer, qubits are still highly susceptible to noise in storage and computation, making it difficult to fulfil the potential of quantum computing. Fortunately, quantum error correction[1, 2] provides a promising approach to deal with this issue, through encoding the logical information into certain subspace of the state space of multiple physical qubits. The choice of the subspace corresponds to a quantum error correcting code (QECC).
@@ I Introduction
Quantum low density parity-check (QLDPC) codes[3] are an important class of QECCs. In 2014, Gotteman proved that making use of QLDPC codes, one can realize fault-tolerant quantum computing with constant overhead[4]. Thus, QLDPC codes have attracted a great deal of attention in recent years. So far, there are many construction methods of QLDPC codes have been proposed, such as hypergraph product[5], homological product[6], lifted product[7], balanced product[8] and XYZ product[9]. All the above construction methods generate CSS codes, except XYZ product which generates non-CSS codes.
CSS codes[10] have some advantages. For instance, decoding a CSS code can be transformed into decoding two classical linear codes, which is much easier than decoding a non-CSS code. However, they are not quite good at dealing with biased noise. The stabilizer generators of a CSS code can be generally divided into two parts: one part only contains Pauli $X$ operators (called $X$-type stabilizer) and the other part only contains Pauli $Z$ operators (called $Z$-type stabilizer). $X$-type stabilizers can only detect Pauli $Z$ and $Y$ errors, while $Z$-type stabilizers can detect Pauli $X$ and $Y$ errors. Due to this, under biased noise, some stabilizers of CSS codes cannot provide information of errors, which may result that the error-correcting performance of a CSS code is unsatisfactory under biased noise. Compared to CSS codes, non-CSS codes perform better under biased noise. We take standard surface code[11, 12] and XZZX surface code[13] for an example. For standard surface code, under pure Pauli $Z$ or $X$ noise, only half of stabilizers can provide error-correcting information for decoder, while under pure Pauli $Y$ noise, all stabilizers can provide error-correcting information for decoder. For XZZX surface code, no matter which noise it’s under, the measurement results of all stabilizer can be utilized by decoder. In other words, they all provide decoder with some information about what the actual error looks like. This is one of the important reasons that under pure $X$ and $Z$ noise the error-correcting performance of XZZX surface code is better than that of standard surface code, while under pure $Y$ noise they have similar error correcting performance. Thus, from this perspective, under biased noise, non-CSS codes tend to have better error-correcting performance compared to CSS codes. In the practical hardware system of quantum computing, the probabilities of Pauli $Z$ noise are much higher than those of Pauli $X$ and $Y$ noise, thus considering quantum error-correcting codes with better performance for biased noise is more meaningful.
@@
XYZ product proposed in Ref. [9] constructs a class of non-CSS codes by using three classical codes, which can be seen as a three-fold variant of the hypergraph product or the standard homological product of three classical codes. However, so far there has been no in-depth study of the error-correcting performance of XYZ product codes, and there are several natural questions: whether this construction method can be generalized to higher dimension? If possible, how is the error-correcting performance of these high-dimensional XYZ product codes, especially under biased noise?
@@
The motivation of our work is to answer the problems above. In this paper, we first study the error-correcting performance of three-dimensional (3D) Chamon codes[14, 15], which is a non-CSS variant of the 3D toric code and a special instance of the XYZ product of three repetition codes with length $n_{1}$, $n_{2}$, $n_{3}$. Exploiting fully decoupled binary belief propagation combined with $order\mbox{-}0$ ordered statistics decoding (FDBP-OSD-0[16]), when $n_{1}=n_{2}=n_{3}$, the code-capacity threshold of the corresponding Chamon codes is around $14\%$ under depolarizing noise model. For pure Pauli $X$, $Y$ and $Z$ noise model, the corresponding code-capacity threshold of Chamon codes is around $13\%$, $14\%$, $14\%$, respectively. Second, we show that that XYZ product can be generalized to four dimension and propose 4D XYZ product code construction, which can be seen as a variant of standard homological product of two CSS codes (namely, 4D homological product) and constructs a non-CSS code by making use of 4 classical codes or 2 CSS codes. Third, we study the dimension and code distance of 4D XYZ product codes, and show that using the same two CSS codes, the non-CSS code constructed by 4D XYZ product may have higher dimension or code distance than those of the CSS code constructed by 4D homological product. Finally, we explore the error-correcting performance of two special instances of 4D XYZ product. The first one is constructed from two quantum concatenated codes, which are obtained from two pairs of repetition codes with block lengths ($n_{1}$, $n_{2}$) and ($n_{3}$, $n_{4}$) (here $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd). The second one is constructed from two hypergraph product codes, which are also obtained from two pairs of repetition codes with block lengths ($n_{1}$, $n_{2}$) and ($n_{3}$, $n_{4}$). We name these two codes 4D XYZ concatenated code and 4D Chamon code, respectively. Exploiting FDBP-OSD-0, we study their error-correcting performances under the independent single-qubit Pauli noise model. Our simulation results show that under pure Pauli $Z$ noise the code-capacity thresholds of 4D XYZ concatenated codes and 4D Chamon codes are much higher than those of the 4D toric codes[17] and 4D homological concatenated codes (namely, the codes constructed by 4D homological product of two quantum concatenated codes, which are the same two codes used to construct 4D XYZ concatenated code) respectively, while their code-capacity thresholds are close under depolarizing noise. Besides, under $Z$-biased noise model with different bias rate, from the perspective of code-capacity threshold, 4D XYZ concatenated codes and 4D Chamon codes also show good error-correcting performance. Our simulation results means that, using the same two codes, 4D XYZ product can constructed non-CSS codes with better error-correcting performance for biased noise than CSS codes constructed by 4D homological product, which is more meaningful for practice quantum computing system.
@@
This paper is organized as follow. Sect. II introduces some preliminaries, including quantum stabilizer codes, chain complex, hypergraph product codes and 3D XYZ product codes. In Sect. III, we introduces the 4D XYZ product code construction and study its dimension and code distance. In Sect. IV, we explore the error-correcting performance of 3D XYZ Chamon codes, 4D Chamon codes and 4D XYZ concatenated codes. In Sect. V, we conclude our work.
II Preliminaries
In this section, we briefly introduce some preliminaries, including quantum stabilizer codes, chain complex, hypergraph product codes and 3D XYZ product codes.
II.1 Quantum stabilizer codes
@@
Quantum stabilizer codes (QSCs)[18] are an important class of quantum error correcting codes, which can be seen as the analogue of classical linear code in the quantum field.
@@
If the stabilizer generators of a QSC $C$ can be divide into two parts, each of which only contains either Pauli $X$ or Pauli $Z$ operators, it is a CSS code, otherwise it is a non-CSS code. It can be seen that each stabilizer generator of a CSS code can only detect two types of Pauli errors, while that of a non-CSS code can detect all three types of Pauli errors.
@@
II.2 Chain complex
In this section, we introduce the concept of chain complex, which will help to better understand hypergraph product code, 3D XYZ product code and our proposed 4D XYZ product code.
@@ II.2 Chain complex
II.3 Hypergraph product and 3D XYZ product codes
@@
Actually, the tensor-product structure depicted in Fig.1 can not only derive a CSS code, but also derive a non-CSS code. As shown in Fig. 1, the stabilizer generators can be divided into two parts, $S$ and $T$, and each part consists of two classes of qubits, $A$ and $B$. In the case of constructing a CSS code, we choose one between $S$ and $T$ to be $X$-type stabilizers and the other to be $Z$-type stabilizers. If we let stabilizer generators in $S$ and $T$ to be mixed $X$ and $Z$ types, we can obtain a non-CSS code. More precisely, let the nontrivial operators of stabilizers in $S$ acting on qubits in $A$ and $B$ be Pauli $X$ and $Z$ operators, respectively. Similarly, let the nontrivial operators of stabilizers in $T$ acting on qubits in $A$ and $B$ be Pauli $Z$ and $X$ operators, respectively. In this way, we obtain the following stabilizer generator matrix,
@@
where the notation $\mathcal{P}=P^{H}\ \left(P\in\left\{X,\ Y,Z\right\}\right)$ denotes a Pauli tensor, which means for any entry of matrix $H$, if it is a 1, $\mathcal{P}$ has a Pauli operator $P$ at the corresponding position, and an identity operator $I$ otherwise. It can be verified any pair of stabilizer generators in $\mathcal{S}$ commutes.
@@
So far, the error-correcting performance of 3D XYZ product codes has not been studied in depth. In Sect. IV, we study the error-correcting performance of 3D Chamon code, which is a non-CSS variant of the 3D toric code and a special instance of the 3D XYZ product of three repetition codes.
@@
III Four-dimensional XYZ product codes
In this section, we first introduce the 4D XYZ product code construction. Second, we give out how to compute the dimension of 4D XYZ product codes and compute two special instances of 4D XYZ product codes. Third, we give out the general form of the logical operators of the 4D XYZ product codes, based on which we prove the upper bound of code distance.
III.1 4D XYZ product code construction
4D XYZ product code construction is using either two CSS codes or four classical linear codes to construct a non-CSS code, since one can use four classical linear codes to construct two CSS codes through hypergraph product or concatenation. Formally, the 4D XYZ product code construction is defined as follow.
@@ III Four-dimensional XYZ product codes
where $m_{1}$, $m_{2}$, $m_{3}$, $m_{4}$, $n_{A}$, $n_{B}$ are the dimensions of $C_{-1}$, $C_{1}$, ${\widetilde{C}}_{-1}$, ${\widetilde{C}}_{1}$, $C_{0}$ and ${\widetilde{C}}_{0}$, respectively. The total number of qubits is
@@
One can see that the tensor-product structure of 4D XYZ product in Fig. 3 is the same as that of 4D homological product, thus it can be seen as a variant of standard 4D homological product of two CSS codes.
@@
To prove the stabilizer group of the 4D XYZ product code is Abelian is to prove any pair of stabilizers commute. Giving the stabilizer matrix $\mathcal{S}$ of the 4D XYZ product code as shown in Eq. (11), the corresponding parity-check matrix $H$ in symplectic representation is $H=\left(H_{x}\mid H_{z}\right)$, where
@@
III.2 Dimension of 4D XYZ product code
In this section, we compute the dimension of the 4D XYZ product code. Similar with Ref. [9], this problem is transformed into finding the number of independent stabilizer generators. First, we give a general solution of this problem in Theorem 1. Second, we consider two special instances of 4D XYZ product codes. The first one is constructed from two concatenated code, which are obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$ (here $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd), and prove that its dimension is 1 in Corollary 2. The second one is constructed from two hypergraph product codes, which are also obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$, and prove that its dimension is $8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$ in Corollary 3. We name these two codes 4D XYZ concatenated code and 4D Chamon code (the reason why we named it this way is explained in Sect. IV.2), respectively.
@@ III.2 Dimension of 4D XYZ product code
where $H_{x_{1}}=\delta_{0}$ and $H_{z_{1}}=\delta_{-1}^{T}$ are the $X$-type and $Z$-type parity-check matrixes of code $C\left(\mathfrak{C}_{1}\right)$, respectively, and $H_{x_{2}}={\widetilde{\delta}}_{0}$ and $H_{z_{2}}={\widetilde{\delta}}_{-1}^{T}$ are the $X$-type and $Z$-type parity-check matrixes of code $C\left(\mathfrak{C}_{2}\right)$, respectively. Let recall that the dimensions of $H_{x_{1}}$, $H_{z_{1}}$, $H_{x_{2}}$ and $H_{z_{2}}$ are $m_{1}\times n_{A}$, $m_{2}\times n_{A}$, $m_{3}\times n_{B}$ and $m_{4}\times n_{B}$, respectively.
In the rest of paper, we use bold lowercase letters to represent column vectors.
Theorem 1 (The dimension of 4D XYZ product code).
The dimension $k$ of 4D XYZ product code is $\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)+k_{SV}+k_{TU}$, where $k_{SV}=\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ is the number of independent solutions $\begin{bmatrix}\textbf{s}\\ \textbf{v}\end{bmatrix}$ of
@@
The total number $m$ of stabilizer generators in Eq. (16) is $m=m_{1}n_{B}+m_{2}n_{B}+m_{3}n_{A}+m_{4}n_{A}$. Thus, according to Eq. (12), the 4D XYZ product code at least encodes $N-m=\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)$ logical qubits. However, we should notice that stabilizer generators in Eq. (16) may not be independent with each other. Thus, we should find out the number of independent stabilizer generators in Eq. (16).
Observing Eq. (16), one can see that the stabilizer generators are divided into $S$, $T$, $U$, $V$ four parts, and any stabilizer in $S$ or $V$ cannot be generated by the product of some stabilizers in $T$ and $U$. Thus, we first consider the independence of the stabilizers in $S$ and $V$, and for the stabilizers in $T$ and $U$ the analysis is the same.
@@
It can be seen that the solution of $I_{m_{1}}\otimes H_{x_{2}}\textbf{\emph{s}}=I_{m_{1}}\otimes H_{z_{2}}\textbf{\emph{v}}=\mathbf{0}$ is $\textbf{\emph{s}}\in\mathcal{C}_{m_{1}}\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ and the solution of $I_{m_{2}}\otimes H_{x_{2}}\textbf{\emph{v}}=I_{m_{2}}\otimes H_{z_{2}}\textbf{\emph{v}}=\mathbf{0}$ is $\textbf{\emph{v}}\in\mathcal{C}_{m_{2}}\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$. Similarly, the solution of $H_{z_{1}}^{T}\otimes I_{n_{B}}\textbf{\emph{s}}=H_{x_{1}}^{T}\otimes I_{n_{B}}\textbf{\emph{v}}$ is $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}\in\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\mathcal{C}_{n_{B}}$. Thus, the solution of equation system (17) is $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}\in\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$, and the dimension of $\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ is the number of independent solutions $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}$ of equation system (17).
Let $k_{SV}=\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$, according to the theory of linear algebra, the number of independent stabilizers $r_{SV}$ is
@@
Similarly, let $k_{TU}=\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)\right)$ is the number of independent solutions $\begin{bmatrix}\textbf{\emph{t}}\\ \textbf{\emph{u}}\end{bmatrix}$, the number of independent stabilizers $r_{TU}$ is
@@
Thus, the total number of independent stabilizers is $r_{SV}+r_{TU}$, and the dimension of 4D XYZ product code is
$N-\left(m_{1}n_{B}+m_{2}n_{B}-k_{SV}\right)-\left(m_{3}n_{A}+m_{4}n_{A}-k_{TU}\right)$ (22)

### v3

@@ Abstract
Three-dimensional (3D) quantum XYZ product can construct a class of non-CSS codes by using three classical codes. However, their error-correcting performance has not been studied in depth so far and whether this code construction can be generalized to higher dimension is an open question. In this paper, we first study the error-correcting performance of the 3D Chamon code, which is an instance of 3D XYZ product of three repetition codes. Next, we show that 3D XYZ product can be generalized to four dimension and propose four-dimensional (4D) XYZ product code construction, which constructs a class of non-CSS codes by using either four classical codes or two CSS codes. Compared with 4D homological product, we show that 4D XYZ product can construct non-CSS codes with higher code dimension or code distance. Finally, we consider two instances of 4D XYZ product, to which we refer as 4D Chamon code and 4D XYZ product concatenated code, respectively. Our simulation results show that, 4D XYZ product can construct non-CSS codes with better error-correcting performance for $Z$-biased noise than CSS codes constructed by 4D homological product.
1 Introduction
@@
Quantum error-correcting code (QECC)[1, 2] provides a promising approach to realize large-scale fault-tolerant quantum computing. In recent years, quantum low density parity-check (QLDPC) codes[3, 4] have attracted a great deal of attention, since they may have high coding rate, large code distance and low stabilizer weight, which benefits their engineering application. So far, many constructions of QLDPC codes have been proposed, such as XYZ product[5], hypergraph product[6], homological product[7], lifted product[8], and balanced product[9]. All the above construction methods generate CSS codes, except XYZ product which generates non-CSS codes.
Compared with CSS codes[10], non-CSS codes have shown better error-correcting performance against biased noise[11, 12]. In quantum computing hardware system, the probabilities of Pauli $Z$ noise are often much higher than those of Pauli $X$ and $Y$ noise[13, 14, 15], thus considering quantum error-correcting codes with better performance against biased noise is more meaningful in this sense.
@@
Three-dimensional (3D) XYZ product proposed in Ref. [5] constructs a class of non-CSS codes by using three classical codes. However, so far there has been no in-depth study of the error-correcting performance of 3D XYZ product codes, and there are several natural questions: whether this construction method can be generalized to higher dimension? If possible, how is the error-correcting performance of these high-dimensional XYZ product codes, especially against biased noise?
@@
The motivation of our work is to answer the problems above. In this paper, we first study the error-correcting performance of 3D Chamon codes[16, 17], which is an instance of 3D XYZ product of three repetition codes with length $n_{1}$, $n_{2}$, $n_{3}$ and is a non-CSS variant of the 3D toric code. Exploiting fully decoupled binary belief propagation combined with $order\mbox{-}0$ ordered statistics decoding (FDBP-OSD-0[18]), when $n_{1}=n_{2}=n_{3}$, the code-capacity threshold of the corresponding Chamon codes is around $14\%$ under depolarizing noise, and around $13\%$, $14\%$, $14\%$ under pure Pauli $X$, $Y$ and $Z$ noise, respectively.
Second, we show that that XYZ product can be generalized to four dimension and propose four-dimensional (4D) XYZ product code construction, which constructs a non-CSS code by making use of either 4 classical codes or 2 CSS codes. It also can be regarded as a non-CSS variant of standard homological product of two CSS codes (namely, 4D homological product).
Third, we show that using the same 2 CSS codes, 4D XYZ product can construct codes with higher code dimension (the number of encoded logical qubits) or code distance than those constructed by 4D homological product.
At last, to explore the error-correcting performance of 4D XYZ product codes, we consider two instances. The first one is constructed from two quantum concatenated codes, which are obtained from two pairs of repetition codes. The second one is constructed from two hypergraph product codes, which are also obtained from two pairs of repetition codes. We refer these two codes to 4D XYZ product concatenated code and 4D Chamon code, respectively. Exploiting FDBP-OSD-0, our simulation results show that under pure Pauli $Z$ noise the code-capacity thresholds of 4D XYZ product concatenated codes and 4D Chamon codes are much higher than those of the 4D toric codes[19] and 4D homological product concatenated codes (namely, the codes constructed by 4D homological product of two quantum concatenated codes, which are the same two codes used to construct 4D XYZ product concatenated code) respectively, while their code-capacity thresholds are close under depolarizing noise. Besides, under $Z$-biased noise model with different bias rate, from the perspective of code-capacity threshold, 4D XYZ product concatenated codes and 4D Chamon codes also show good error-correcting performance. Our simulation results means that, using the same two codes, 4D XYZ product can constructed non-CSS codes with better error-correcting performance against biased noise than CSS codes constructed by 4D homological product.
@@
This paper is organized as follow. Sect. 2 introduces some preliminaries, including quantum stabilizer codes, chain complex, hypergraph product and 3D XYZ product code construction. In Sect. 3, we introduce the 4D XYZ product code construction and study its code dimension and code distance. In Sect. 4, we explore the error-correcting performance of 3D XYZ Chamon codes, 4D Chamon codes and 4D XYZ product concatenated codes. In Sect. 5, we conclude our work.
2 Preliminaries
2.1 Quantum stabilizer codes
@@
Quantum stabilizer codes (QSCs)[20] are an important class of quantum error correcting codes, which can be seen as the analogue of classical linear codes in the quantum field.
@@
If the stabilizer generators of a QSC $C$ can be divided into two parts, each of which only contains either Pauli $X$ or Pauli $Z$ operators, it is a CSS code, otherwise it is a non-CSS code. It can be seen that each stabilizer generator of a CSS code can only detect two types of Pauli errors, while that of a non-CSS code can detect all three types of Pauli errors.
@@
2.2 Chain complex
In this section, we introduce the concept of chain complex, which will help to better understand the hypergraph product code, 3D XYZ product code and our proposed 4D XYZ product code.
@@ 2.2 Chain complex
2.3 Hypergraph product and 3D XYZ product code construction
@@
The tensor-product structure depicted in Fig. 1 can derive not only a CSS code but also a non-CSS code. As shown in Fig. 1, the stabilizer generators can be divided into two parts, $S$ and $T$, and each part consists of two classes of qubits, $A$ and $B$. In the case of constructing a CSS code, we assign $X$-type stabilizers to one of $S$ or $T$ and $Z$-type stabilizers to the other. If we let stabilizer generators in $S$ and $T$ to be mixed $X$ and $Z$ types, we can obtain a non-CSS code. More precisely, let the nontrivial operators of stabilizers in $S$ acting on qubits in $A$ and $B$ be Pauli $X$ and $Z$ operators, respectively. Similarly, let the nontrivial operators of stabilizers in $T$ acting on qubits in $A$ and $B$ be Pauli $Z$ and $X$ operators, respectively. In this way, we obtain the following stabilizer generator matrix,
@@
where the notation $\mathcal{P}=P^{H}\ \left(P\in\left\{X,\ Y,Z\right\}\right)$ denotes a Pauli tensor, which means for any entry of matrix $H$, if the entry is 1, the tensor $\mathcal{P}$ places a Pauli operator $P$ at the corresponding position, and an identity operator $I$ otherwise. It can be verified any pair of stabilizer generators in $\mathcal{S}$ commutes.
@@
So far, the error-correcting performance of 3D XYZ product codes has not been studied in depth. In Sect. 4, we study the error-correcting performance of 3D Chamon code, which is a non-CSS variant of the 3D toric code and an instance of the 3D XYZ product of three repetition codes.
@@
3 Four-dimensional XYZ product codes
In this section, we first introduce the 4D XYZ product code construction. Second, we explain how to compute the code dimension of 4D XYZ product codes and compute two instances’ code dimension. Third, we give out the general form of the logical operators of the 4D XYZ product codes, based on which we prove the upper bound of code distance.
3.1 4D XYZ product code construction
The construction of 4D XYZ product code involves either two CSS codes or four classical linear codes to create a non-CSS code, since one can use four classical codes to construct two CSS codes through hypergraph product or concatenation. Formally, the 4D XYZ product code construction is defined as follows.
@@ 3 Four-dimensional XYZ product codes
where $m_{1}$, $m_{2}$, $m_{3}$, $m_{4}$, $n_{A}$, $n_{B}$ are the dimensions of spaces $C_{-1}$, $C_{1}$, ${\widetilde{C}}_{-1}$, ${\widetilde{C}}_{1}$, $C_{0}$ and ${\widetilde{C}}_{0}$, respectively. The total number of qubits is
@@
One can see that the tensor-product structure of 4D XYZ product in Fig. 3 is the same as that of 4D homological product. Therefore, it can be considered as a variant of the standard 4D homological product of two CSS codes.
@@
To prove that the stabilizer group of the 4D XYZ product code is Abelian, we need to prove that any pair of stabilizers commute. Given the stabilizer matrix $\mathcal{S}$ of the 4D XYZ product code as shown in Eq. (11), the corresponding parity-check matrix $H$ in symplectic representation is $H=\left(H_{x}\mid H_{z}\right)$, where
@@
3.2 Code dimension of 4D XYZ product code
In this section, we compute the code dimension of the 4D XYZ product codes. Similar to Ref. [5], this problem reduces to finding the number of independent stabilizer generators. First, we give a general solution to this problem in Theorem 1. Second, we consider two instances of 4D XYZ product. The first one is constructed from two concatenated codes, which are obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$ (here $n_{1}$, $n_{2}$, $n_{3}$ and $n_{4}$ are all odd), and its code dimension is proven to be 1 in in Corollary 2. The second one is constructed from two hypergraph product codes, which are also obtained from two pairs of repetition codes with block lengths $\left(n_{1},\ n_{2}\right)$ and $\left(n_{3},\ n_{4}\right)$, and its code dimension is proven to be $8\gcd(n_{1},n_{2})\gcd(n_{3},n_{4})$ in Corollary 3. We refer to these two codes as the 4D XYZ concatenated code and the 4D Chamon code (the rationale for this name is discussed in Sect. 4.2), respectively.
@@ 3.2 Code dimension of 4D XYZ product code
where $H_{x_{1}}=\delta_{0}$ ($H_{x_{2}}={\widetilde{\delta}}_{0}$) and $H_{z_{1}}=\delta_{-1}^{T}$ ($H_{z_{2}}={\widetilde{\delta}}_{-1}^{T}$) are the $X$-type and $Z$-type parity-check matrices of code $C\left(\mathfrak{C}_{1}\right)$ [$C\left(\mathfrak{C}_{2}\right)$], respectively. Recall that the dimensions of $H_{x_{1}}$, $H_{z_{1}}$, $H_{x_{2}}$ and $H_{z_{2}}$ are $m_{1}\times n_{A}$, $m_{2}\times n_{A}$, $m_{3}\times n_{B}$ and $m_{4}\times n_{B}$, respectively.
For the remainder of this paper, bold lowercase letters represent column vectors.
Theorem 1 (The code dimension of 4D XYZ product code).
The code dimension $k$ of 4D XYZ product code is $\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)+k_{SV}+k_{TU}$, where $k_{SV}=\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ is the number of independent solutions $\begin{bmatrix}\textbf{s}\\ \textbf{v}\end{bmatrix}$ of
@@
The total number $m$ of stabilizer generators in Eq. (16) is $m=m_{1}n_{B}+m_{2}n_{B}+m_{3}n_{A}+m_{4}n_{A}$. Thus, according to Eq. (12), the 4D XYZ product code encodes at least $N-m=\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)$ logical qubits. However, we should notice that stabilizer generators in Eq. (16) may not be independent of each other. Thus, we should find out the number of independent stabilizer generators in Eq. (16).
Observing Eq. (16), one can see that the stabilizer generators are divided into $S$, $T$, $U$, $V$ four parts, and any stabilizer in $S$ or $V$ cannot be generated by the product of some stabilizers in $T$ and $U$. Thus, we first consider the independence of the stabilizers in $S$ and $V$, with a similar analysis applied to the stabilizers in T and U.
@@
It can be seen that the solution of $I_{m_{1}}\otimes H_{x_{2}}\textbf{\emph{s}}=I_{m_{1}}\otimes H_{z_{2}}\textbf{\emph{s}}=\mathbf{0}$ is $\textbf{\emph{s}}\in\mathcal{C}_{m_{1}}\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ and the solution of $I_{m_{2}}\otimes H_{x_{2}}\textbf{\emph{v}}=I_{m_{2}}\otimes H_{z_{2}}\textbf{\emph{v}}=\mathbf{0}$ is $\textbf{\emph{v}}\in\mathcal{C}_{m_{2}}\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$. Similarly, the solution of $H_{z_{1}}^{T}\otimes I_{n_{B}}\textbf{\emph{s}}=H_{x_{1}}^{T}\otimes I_{n_{B}}\textbf{\emph{v}}$ is $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}\in\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\mathcal{C}_{n_{B}}$. Thus, the solution of equation system (17) is $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}\in\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$, and the dimension of $\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ is the number of independent solutions $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}$ of equation system (17).
Notice that
$k_{SV}=\dim\left(\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ (20)
and according to the theory of linear algebra, the number of independent stabilizers in $S$ and $V$, $r_{SV}$, is
@@
Similarly, notice that $k_{TU}=\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)\right)$ is the number of independent solutions $\begin{bmatrix}\textbf{\emph{t}}\\ \textbf{\emph{u}}\end{bmatrix}$ of equation system (18), thus the number of independent stabilizers in $T$ and $U$, $r_{TU}$, is
@@
Thus, the total number of independent stabilizers in Eq. (16) is $r_{SV}+r_{TU}$, and the code dimension of 4D XYZ product code is
$N-r_{SV}-r_{TU}$ (23)